# The CLAS Collaboration

### v2

@@ VIII Extracted quasi-free cross sections
In Fig. 14, the $W$ dependences of the extracted fully integrated cross sections of the reaction $\gamma_{v}p(n)\rightarrow p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$ are shown by the black filled circles for twelve analyzed $Q^{2}$ bins. For each point, the pink shadowed area is the total cross section uncertainty, which is the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ (see Sec. VI.4) summed up in quadrature with the total systematic uncertainty (see Sec. VI.3). The error bars correspond to the $\delta_{\text{stat,mod}}^{\text{tot}}$ uncertainty only.
[…]
@@
The extracted cross sections are quasi-free, meaning that the contribution from events in which the final hadrons interacted with the spectator neutron is reduced to the kinematically achievable minimum. The cross sections, however, are still convoluted with effects of FSIs among the final hadrons, which is like conventional free proton cross sections (see Sec. VII.2). Also note that in this study, the initial proton is assumed to be on-shell.
@@
In general, the admixture of the FSI-background left after the exclusivity cut in the $\pi^{-}$ missing topology may potentially affect the shape of the extracted single-differential distributions as it was corrected only in an integral sense (as described in Sec. III.5.4). However, as this admixture is present only for events from the $\pi^{-}$ missing topology for $W>$ 1.45 GeV and stays there at a level of 3%-7%, its impact is not thought to be discernible against the total cross section uncertainty.
@@
Figure 16: Ratio of the fully integrated quasi-free cross sections obtained in this study over the free proton cross sections from Refs. Fedotov et al. (2017, 2018). The red curves correspond to the polynomial fit. The dashed line marks the value of 0.75.
@@
One more potential uncertainty source for the extracted quasi-free cross sections is the presence of falsely identified quasi-free events in the $\pi^{-}$ missing topology. Unfortunately, as true quasi-free events are kinematically identical to those that are falsely identified, no corresponding correction to the cross section can be developed (see Sec. VII.5 for details).
@@
The solid curves in Figs. 14 and 15 correspond to the cross section estimation performed by means of TWOPEG Skorodumina et al. (2017b), which is the event generator for double-pion electroproduction off free protons. This event generator currently provides the best estimation of the free proton cross sections in the investigated kinematic region.
@@
The cross section approximation implemented into the TWOPEG event generator is based on the meson-baryon reaction model JM Mokeev et al. (2009, 2012, 2016). The generator employs the five-differential structure functions from the recent version of the JM model fit to the existing CLAS results on double-pion photo- and electroproduction off free protons Ripani et al. (2003); Mokeev et al. (2012); Fedotov et al. (2009); Golovatch et al. (2019). In the kinematic areas already covered by the CLAS data, TWOPEG performs the interpolation of the model structure functions and successfully reproduces the available fully integrated and single-differential cross sections. In the areas not yet covered by the CLAS data, special extrapolation procedures were applied that included additional world data on the fully integrated photoproduction cross sections Wu et al. (2005); ABBHHM Collaboration (1968).
@@
For the purpose of this comparison, the cross section distributions obtained by TWOPEG were normalized to integrally match the quasi-free cross sections extracted in this study.
@@
Figures 14 and 15 indicate that apart from the overall integral scaling, the free proton cross sections may serve as an adequate zeroth-order approximation for the quasi-free cross sections off protons in deuterium. More elaborate insight into the interpretation of the extracted quasi-free cross sections is given in the next section, which presents their comparison with the free proton measurements from Refs. Fedotov et al. (2017, 2018).
@@ IX Comparison with the free proton measurements
Figure 17: Ratio of the invariant mass distributions obtained in this study over their free proton analogues from Refs. Fedotov et al. (2017, 2018). Rows from top to bottom correspond to $M_{p^{\prime}\pi^{+}}$, $M_{\pi^{+}\pi^{-}}$, and $M_{\pi^{-}p^{\prime}}$, respectively. The ratios were averaged over $Q^{2}$ to decrease the resulting uncertainties. The red curves correspond to polynomial fits. The dashed line marks the value of 0.75.
@@
Figure 18: Ratio of the angular distributions from this study over their free proton analogues from Refs. Fedotov et al. (2017, 2018). The first row shows the $\theta_{p^{\prime}}$, $\theta_{\pi^{-}}$, and $\theta_{\pi^{+}}$ distributions, while the second row shows the $\alpha_{p^{\prime}}$, $\alpha_{\pi{-}}$, and $\alpha_{\pi{+}}$ distributions, respectively. The ratios were averaged over $W$ and $Q^{2}$ to minimize the resulting uncertainties. The red curves correspond to polynomial fits. The dashed line marks the value of 0.75.
@@
This study benefits from the fact that the cross sections of the same exclusive reaction off free protons have been recently extracted from CLAS data Fedotov et al. (2017, 2018). These free proton measurements were performed under the same experimental conditions as in this study, including the beam energy value and the target setup. For most points of the fully integrated cross sections, the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ is at a level of 1%-3% for the free proton measurements and at a level of 4%-6% for the cross sections obtained in this study (see Sec. VI.4). The two measurements have identical binning in all kinematic variables and similar inherent systematic inaccuracies. A typical value of the total systematic uncertainty in a $W$ and $Q^{2}$ bin is about 7%-8% for both cross section sets. Therefore, a direct comparison of the cross sections extracted in this study with their free proton counterparts from Refs. Fedotov et al. (2017, 2018) provides the experimentally best possible opportunity to explore distinctions between the $\pi^{+}\pi^{-}$ electroproduction off protons in deuterium and the corresponding reaction off free protons.
@@
This section compares the two cross section sets and examines the difference between them over the entire reaction phase space. In this investigation, only the statistical and model-dependent uncertainties of the two measurements were considered ($\delta_{\text{stat,mod}}^{\text{tot}}$), while the systematic effects were assumed to cancel out.
@@
Figure 16 presents the ratio of the fully integrated quasi-free cross sections obtained in this study over the free proton cross sections of Refs. Fedotov et al. (2017, 2018). The ratio was fit with a fourth-order polynomial. The dashed line marks the value of 0.75.
@@
As seen in Fig. 16, the ratio of the two cross section sets demonstrates a modest $W$ dependence with an average level of 70%-75%, appearing to drop slightly near the threshold, as well as in the dip region between the two integral resonance peaks.
@@
Figure 17 shows the ratio of the invariant mass distributions from this study and the free proton study Fedotov et al. (2017, 2018). Rows from top to bottom correspond to $M_{p^{\prime}\pi^{+}}$, $M_{\pi^{+}\pi^{-}}$, and $M_{\pi^{-}p^{\prime}}$, respectively. The ratios were obtained individually for each $(W,Q^{2})$ point and then averaged over $Q^{2}$ to decrease the resulting uncertainties. The kinematic broadening of the invariant mass distributions with $W$ (illustrated by Eq. (8) in Sec. IV.4) and the consequent nonidentical distribution of data points in different $W$ bins does not allow for further averaging over $W$. The red curves correspond to the fit with a fourth-order polynomial and the dashed line marks the value of 0.75.
@@
As seen in Fig. 17, the cross section ratio demonstrates different consistent patterns for the three invariant mass distributions. For $M_{p^{\prime}\pi^{+}}$ (top row), it gives a rise near the left distribution edge, then gradually drops towards the right edge, featuring a small plateau in the middle. For $M_{\pi^{+}\pi^{-}}$ (middle row), the situation is different, i.e. the cross section ratio shows a pronounced drop of up to $\sim$40at the left edge, then rises abruptly up to $\sim$75and stays on this constant level further on. For the third invariant mass, $M_{\pi^{-}p^{\prime}}$ (bottom row), the ratio continuously and almost linearly grows from $\sim$60at the left edge to $\sim$100at the right.
@@
Figure 18 presents the ratio of the angular distributions from this study and the free proton study Fedotov et al. (2017, 2018). The first row shows the $\theta_{p^{\prime}}$, $\theta_{\pi^{-}}$, and $\theta_{\pi^{+}}$ distributions, while the second row shows the $\alpha_{p^{\prime}}$, $\alpha_{\pi{-}}$, and $\alpha_{\pi{+}}$ distributions, respectively. The ratios were obtained individually for each $(W,Q^{2})$ point and then averaged over $W$ and $Q^{2}$ to minimize the resulting uncertainties. The red curves correspond to polynomial fits and the dashed line marks the value of 0.75.
@@
As seen in Fig. 18, the behavior of the cross section ratio differs for various angular distributions. Its dependence on hadron polar angles appears to be of the most interest. For $\theta_{p^{\prime}}$ (top left plot), the ratio starts from $\sim$70at small angles, grows up to $\sim$75at ${}^{\circ}$, and further stays on a distinct plateau up to ${}^{\circ}$, showing then a mild rise up to $\sim$80at backward angles. For $\theta_{\pi^{-}}$ (top middle plot), the ratio stays at $\sim$75at small angles, then grows up to $\sim$80giving a broad peak at around ${}^{\circ}$, and drops down to $\sim$70at backward angles. For the third polar angle ($\theta_{\pi^{+}}$), the ratio maintains the level of $\sim$75all the way through around ${}^{\circ}$, then peaks up to more than 80% at ${}^{\circ}$, and finally shows a steep drop.
@@
Note that the conventional free proton cross sections represent all reaction events, while the cross sections extracted in this study are quasi-free (up to the accuracy with which quasi-free events can be kinematically isolated) and hence do not include contributions from events in which final hadrons interacted with the spectator neutron. The latter events therefore are mainly responsible for the difference between the two cross section sets.
@@
Therefore, the performed comparison allows us to estimate the proportion of events affected by FSIs with the spectator neutron for the reaction off protons in deuterium. From Figs. 1617, and 18, one can conclude that the contribution from such events to the total number of reaction events varies from $\sim$60to a few percent in different regions of the reaction phase space. However, for the most part of the phase space, one can estimate the contribution from events affected by FSIs with the spectator to be on a level of $\sim$25%.
@@ X CONCLUSIONS
[…]
Due to the Fermi motion that initial protons undergo in deuterium nuclei, this study encountered a set of peculiarities, which were not relevant for free proton studies. Effects of the initial proton motion turned out to be intertwined with many analysis aspects: they lead to the smearing of some kinematic quantities, alter the common procedure of the Lab to CMS transformation, cause the need to perform an unfolding correction to the extracted cross sections, and more Skorodumina et al. (2020); Skorodumina ; Skorodumina et al. (2017a). To deal with these issues, special methods and techniques were developed, which go beyond the conventional analysis framework elaborated in previous free proton studies Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Golovatch et al. (2019); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018).
Interactions of the reaction final hadrons with the spectator nucleon represent another peculiar aspect of this analysis. These interactions introduce distortions into distributions of some kinematic quantities (such as missing masses) thus complicating the exclusive event selection. FSI effects have been found to differ depending on the reaction topology due to nonidentical geometrical acceptance of the topologies in CLAS. For this study, isolation of quasi-free events from the FSI-background was performed in each topology individually according to the specially developed procedures.
The paper also presents the comparison of the obtained cross sections with the corresponding free proton cross sections recently extracted from CLAS data Fedotov et al. (2017, 2018). Assuming that the difference between the two cross section sets mostly originates from events in which FSIs between the final hadrons and the spectator neutron took place, this comparison allowed us to make an estimate of the contribution from such events to the total number of reaction events. For the most part of the reaction phase space, the contribution from events affected by FSIs with the spectator was found to be around 25%. Further theoretical interpretation of the obtained cross section ratios will allow one to investigate different involved FSI mechanisms and also to explore other potential reasons that may contribute to the difference between the cross section sets, which includes possible in-medium modifications of properties of nucleons and their excited states Mokeev et al. (1995); Bianchi et al. (1994); Ahrens (1985); Krusche (2005); Noble (1981).

### v4

@@ VIII Extracted quasifree cross sections
In Fig. 14, the $W$ dependences of the extracted fully integrated cross sections of the reaction $\gamma_{v}p(n)\rightarrow p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$ are shown by the solid black circles for 12 analyzed $Q^{2}$ bins. For each point, the pink shadowed area is the total cross section uncertainty, which is the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ (see Sec. VI.4) summed up in quadrature with the total systematic uncertainty (see Sec. VI.3). The error bars correspond to the $\delta_{\text{stat,mod}}^{\text{tot}}$ uncertainty only.
For each integral cross section point, a set of nine single-differential cross sections was obtained (as described in Sec. IV.5). As a typical example, Figure 15 presents the single-differential cross sections for $W=1.5375$ GeV and $Q^{2}=0.625$ GeV2. The cross sections are reported with the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ shown by the error bars. The full set of extracted single-differential cross sections is available in the CLAS physics database CLAS Physics Database and also on GitHub Git .
@@
The extracted cross sections are quasifree, meaning that the contribution from events in which the final hadrons interacted with the spectator neutron was reduced to the kinematically achievable minimum. The cross sections, however, are still convoluted with effects of FSIs among the final hadrons, which is like conventional free proton cross sections (see Sec. VII.2).
@@
In general, the admixture of the FSI background left after the exclusivity cut in the $\pi^{-}$ missing topology may potentially affect the shape of the extracted single-differential distributions (because this effect was corrected only in an integral sense as described in Sec. III.5.4). However, as this admixture is present only in the $\pi^{-}$ missing topology for $W>$ 1.45 GeV and stays at a level of 3-7%, its impact is not thought to be discernible against the total cross section uncertainty.
@@
Figure 16: Ratio of the fully integrated quasifree cross sections obtained in this study to the free proton cross sections from Refs. Fedotov et al. (2017, 2018). The red curves correspond to polynomial fits. The dashed line marks the value of 0.75.
@@
One more potential uncertainty source for the extracted quasifree cross sections is the presence of falsely identified quasifree events in the $\pi^{-}$ missing topology. Unfortunately, as true quasifree events are kinematically identical to those that are falsely identified, no corresponding correction to the cross sections can be developed (see Sec. VII.5 for details). Also note that in this study, the initial proton is assumed to be on shell.
@@
The solid curves in Figs. 14 and 15 correspond to the cross section estimation performed by means of TWOPEG Skorodumina et al. (2017b), which is an event generator for double-pion electroproduction off free protons. This event generator currently provides the best estimation of free proton cross sections in the investigated kinematic region.
@@
The cross section approximation implemented into the TWOPEG event generator is based on the meson-baryon reaction model JM Mokeev et al. (2009, 2012, 2016). The generator employs the five-differential structure functions from the JM model fit to the existing CLAS results on double-pion photo- and electroproduction off free protons Ripani et al. (2003); Mokeev et al. (2012); Fedotov et al. (2009); Golovatch et al. (2019). In the kinematic areas already covered by the CLAS data, TWOPEG performs interpolation of the model structure functions and successfully reproduces the available fully integrated and single-differential cross sections. In the areas not yet covered by the CLAS data, special extrapolation procedures were applied that included additional world data on the fully integrated photoproduction cross sections Wu et al. (2005); ABBHHM Collaboration (1968).
@@
For the comparison shown in Figs. 14 and 15, cross section distributions obtained by TWOPEG were normalized to integrally match the quasifree cross sections extracted in this study.
@@
Figures 14 and 15 indicate that apart from the overall integral scaling, free proton cross sections may serve as an adequate zeroth-order approximation for quasifree cross sections off protons in deuterium. More elaborate insight into the interpretation of the extracted cross sections is given in the next section, which presents their comparison with the free proton measurements from Refs. Fedotov et al. (2017, 2018).
@@ IX Comparison with the free proton measurements
Figure 17: Ratio of the invariant mass distributions obtained in this study to their free proton counterparts from Refs. Fedotov et al. (2017, 2018). Rows from top to bottom correspond to $M_{p^{\prime}\pi^{+}}$, $M_{\pi^{-}\pi^{+}}$, and $M_{\pi^{-}p^{\prime}}$, respectively. The ratios were averaged over $Q^{2}$ to decrease the resulting uncertainties. The red curves correspond to polynomial fits. The dashed line marks the value of 0.75.
@@
Figure 18: Ratio of the angular distributions from this study to their free proton counterparts from Refs. Fedotov et al. (2017, 2018). The first row shows the $\theta_{p^{\prime}}$, $\theta_{\pi^{-}}$, and $\theta_{\pi^{+}}$ distributions, while the second row shows the $\alpha_{p^{\prime}}$, $\alpha_{\pi^{-}}$, and $\alpha_{\pi^{+}}$ distributions, respectively. The ratios were averaged over $W$ and $Q^{2}$ to minimize the resulting uncertainties. The red curves correspond to polynomial fits. The dashed line marks the value of 0.75.
@@
This study benefits from the fact that the cross sections of the same exclusive reaction off free protons have been recently extracted from CLAS data Fedotov et al. (2017, 2018). These free proton measurements were performed under the same experimental conditions as in this study, including the beam energy value and the target setup. For most points of the fully integrated cross sections, the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ is at a level of 1-3% for the free proton measurements and at a level of 4-6% for the cross sections obtained in this study (see Sec. VI.4). The two measurements have identical binning in all kinematic variables and similar inherent systematic inaccuracies (with a typical value of the total systematic uncertainty in a $W$ and $Q^{2}$ bin of about 7-8%). Therefore, a direct comparison of these two cross section sets would provide the experimentally best possible opportunity to explore distinctions between the $\pi^{+}\pi^{-}$ electroproduction off protons in deuterium and the corresponding reaction off free protons.
@@
This section compares the cross sections extracted in this study with their free proton counterparts from Refs. Fedotov et al. (2017, 2018) and examines the difference between them over the entire reaction phase space. In this comparison, only statistical and model-dependent uncertainties of the two measurements were considered ($\delta_{\text{stat,mod}}^{\text{tot}}$), while systematic effects were assumed to cancel out.
@@
Figure 16 presents the ratio of the fully integrated quasifree cross sections obtained in this study to the free proton cross sections of Refs. Fedotov et al. (2017, 2018). The ratio was fit with a fourth-order polynomial. The dashed line marks the value of 0.75.
@@
As seen in Fig. 16, the ratio of the two cross section sets demonstrates a modest $W$ dependence with an average level of 70-75%, appearing to drop slightly near the reaction threshold, as well as in the dip region between the two integral resonance peaks.
@@
Figure 17 shows the ratio of the invariant mass distributions from this study to those from the free proton study Fedotov et al. (2017, 2018). Rows from top to bottom correspond to $M_{p^{\prime}\pi^{+}}$, $M_{\pi^{-}\pi^{+}}$, and $M_{\pi^{-}p^{\prime}}$, respectively. The ratios were obtained individually for each $(W,Q^{2})$ point and then averaged over $Q^{2}$ to decrease the resulting uncertainties. Kinematic broadening of the invariant mass distributions with $W$ (described in Sec. IV.4) and the consequent nonidentical distribution of data points in different $W$ bins did not allow for further averaging over $W$. The red curves correspond to the fit with a fourth-order polynomial and the dashed line marks the value of 0.75.
@@
As seen in Fig. 17, the cross section ratio demonstrates different consistent patterns for the three invariant mass distributions. For $M_{p^{\prime}\pi^{+}}$ (top row), it gives a rise near the left distribution edge, then gradually decreases towards the right edge, featuring a small plateau in the middle. For $M_{\pi^{-}\pi^{+}}$ (middle row), the situation is different as the cross section ratio shows a pronounced drop down to $\approx$40at the left edge, then rises abruptly up to $\approx$75and stays on this constant level further on. For the third invariant mass, $M_{\pi^{-}p^{\prime}}$ (bottom row), the ratio continuously and almost linearly grows from $\approx$60at the left edge to $\approx$100at the right.
@@
Figure 18 presents the ratio of the angular distributions from this study to those from the free proton study Fedotov et al. (2017, 2018). The first row shows the $\theta_{p^{\prime}}$, $\theta_{\pi^{-}}$, and $\theta_{\pi^{+}}$ distributions, while the second row shows the $\alpha_{p^{\prime}}$, $\alpha_{\pi^{-}}$, and $\alpha_{\pi^{+}}$ distributions, respectively. The ratios were obtained individually for each $(W,Q^{2})$ point and then averaged over $W$ and $Q^{2}$ to minimize the resulting uncertainties. The red curves correspond to polynomial fits and the dashed line marks the value of 0.75.
@@
As seen in Fig. 18, the behavior of the cross section ratio differs for various angular distributions. Its dependence on the hadron polar angles appears to be of the most interest. For $\theta_{p^{\prime}}$ (top left plot), the ratio grows from $\approx$70at small angles up to $\approx$75at 50 and further stays on a distinct plateau up to 120, showing then a mild rise up to $\approx$80at backward angles. For $\theta_{\pi^{-}}$ (top middle plot), the ratio stays at $\approx$75at small angles, then grows up to $\approx$80giving a broad peak at around 100, and drops down to $\approx$70at backward angles. For the third polar angle, $\theta_{\pi^{+}}$ (top right plot), the ratio maintains the level of $\approx$75up to around 120, then peaks up to more than 80% at 150, and finally shows a steep drop.
@@
Note that conventional free proton cross sections represent all reaction events, while the cross sections extracted in this study are quasifree (up to the accuracy with which quasifree events could be kinematically isolated) and hence do not include contributions from events in which the final hadrons interacted with the spectator neutron. The latter events thus are mainly responsible for the difference between the two cross section sets.
@@
Therefore, the performed comparison allows us to estimate the proportion of events affected by FSIs with the spectator neutron for the reaction off protons in deuterium. From Figs. 1617, and 18, one can conclude that the contribution from such events to the total number of reaction events varies from $\approx$60to a few percent in different regions of the reaction phase space. However, for the most part of the phase space, one can estimate the contribution from events affected by FSIs with the spectator to be on a level of $\approx$25%.
@@ X CONCLUSIONS
The fully integrated and single-differential cross sections of the reaction $\gamma_{v}p(n)\rightarrow p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$ have been obtained for the first time. The measurements were performed in the kinematic region of the invariant mass $W$ from 1.3 to 1.825 GeV and the photon virtuality $Q^{2}$ from 0.4 GeV to 1.0 GeV2. The results benefit from fine binning in all kinematic variables, small statistical uncertainties, and modest model dependence. The extracted cross sections are quasifree, meaning that the admixture of events in which the final hadrons interacted with the spectator neutron is reduced to the kinematically achievable minimum. The whole set of the obtained cross sections is available in the CLAS physics database CLAS Physics Database and also on GitHub Git .
Due to the Fermi motion that initial protons undergo in deuterium nuclei, this study encountered a set of peculiarities that were not relevant for free proton studies. Effects of the initial proton motion turned out to be intertwined with many analysis aspects: they led to the smearing of some kinematic quantities, altered the common procedure of the Lab-to-CMS transformation, caused the need to perform an unfolding correction to the extracted cross sections, and more Skorodumina et al. (2020); Skorodumina ; Skorodumina et al. (2017a, 2021). To deal with these issues, special methods and techniques were developed, which go beyond the conventional analysis framework elaborated in previous free proton studies Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Golovatch et al. (2019); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018).
Interactions of the reaction final hadrons with the spectator nucleon represent another peculiar aspect of this analysis. These interactions introduce distortions into distributions of some kinematic quantities (such as missing masses), thus complicating the exclusive event selection. FSI effects were found to differ depending on the reaction topology due to nonidentical geometrical acceptance of the topologies in CLAS. For this study, isolation of quasifree events from the FSI background was performed in each topology individually according to the specially developed procedures.
The paper also presents the comparison of the obtained cross sections with the corresponding free proton cross sections recently extracted from CLAS data Fedotov et al. (2017, 2018). Assuming that the difference between the two cross section sets mostly originates from events in which FSIs between the final hadrons and the spectator neutron took place, this comparison allowed us to make an estimate of the contribution from such events to the total number of reaction events. For the most part of the reaction phase space, the contribution from events affected by FSIs with the spectator was found to be around 25%. Further theoretical interpretation of the obtained cross section ratios will allow one to investigate various involved FSI mechanisms and to explore other potential reasons that may contribute to the difference between the cross section sets, which includes possible in-medium modifications of properties of nucleons and their excited states Mokeev et al. (1995); Bianchi et al. (1994); Ahrens (1985); Krusche (2005); Noble (1981).